OpenStudy (anonymous):

A particle along a straight line with uniform acceleration covers distance AB= 8m and BC=18m in one second each . The speed of the paticle at point B is?

OpenStudy (anonymous):



OpenStudy (anonymous):

The answer should be 13 m/s

OpenStudy (ajprincess):

Use s=ut+1/2at^2

OpenStudy (ajprincess):

First for AB and then for AC. Solve them to find u and a. Then using these infos find the velocity.

OpenStudy (anonymous):

@ajprincess both acceleration and u is not given so how???

OpenStudy (ajprincess):

u hav to find them.

OpenStudy (anonymous):

Let me try and will inform u ....

OpenStudy (ajprincess):

ha k.:)

OpenStudy (anonymous):

@ajprincess After solving i got a=10m/s^2

OpenStudy (ajprincess):

that's right. nw find u.

OpenStudy (anonymous):

yes u=3m/s

OpenStudy (anonymous):

so 13m/s

OpenStudy (anonymous):

thxxx for helping

OpenStudy (ajprincess):

yw. well done.:)

OpenStudy (anonymous):

\[s=ut+1/2 at ^{2}\] \[8=u _{1}+a/2--- eq(1)\] \[18=u _{2}+a/2---eq(2)\] dividing eq(1) and eq(2) \[\left[ (8-u _{1})/18-u _{2} \right]=1\] \[8-u _{1}=18-u _{2}\] \[u _{2}-u _{1}=10\] as\[u _{1}=3\] \[u _{2}=13 m/s\]